# MASA: Rethinking the Representational Bottleneck in LoRA with Multi-A Shared Adaptation

### v1

@@ Abstract
Low-Rank Adaptation (LoRA) has emerged as a dominant method in Parameter-Efficient Fine-Tuning (PEFT) for large language models, which augments the transformer layer with one down-projection $A$ and one up-projection $B$. However, LoRA’s reliance on a single down-projection matrix ($A$) creates a representational bottleneck, as this solitary feature extractor is inherently insufficient for capturing the diverse signals required by complex tasks. This motivates our architectural shift to focus on enriching the feature adaptation to improve the downstream task adaptation ability. We propose MASA (Multi-$A$ Shared Adaptation), an architecture that implements a “multi-$A$, single-BB” structure where the multi-$A$ expert ensemble is asymmetrically shared across layers to ensure parameter efficiency. In MASA, these specialized experts capture diverse features, which are then integrated by a single, layer-specific $B$-matrix. The effectiveness and versatility of our method are validated through a comprehensive suite of experiments spanning multi-domain generalization, single-domain specialization, and multi-task reasoning. For example, on the MMLU benchmark, MASA achieves an average accuracy of 59.62%, outperforming the standard LoRA by 1.08 points (a relative improvement of 1.84%) with comparable learnable parameters of 0.52%.
Introduction
@@
Figure 1: Architectures of LoRA variants and our proposed MASA. (a) LoRA: Each fine‑tuned layer is augmented with a single pair of low‑rank adapters, one down‑projection AA and one up‑projection $B$. (b) LoRA+MoE: Model capacity is increased by instantiating $k$ independent adapter pairs $(A_{i},B_{i})$. (c) HydraLoRA: Several up‑projection heads $B_{i}$ share a common down‑projection AA, forming a “single‑AA, multi‑BB” topology for parameter reuse. (d) Ours: We employ one shared up‑projection $B$ with multiple down‑projections $A_{i}$, which offers a balanced trade‑off between efficiency and representational capacity.
@@
Large Language Models (LLMs) (Brown et al. 2020; Devlin et al. 2019; Touvron et al. 2023; Aydın and Karaarslan 2022; OpenAI 2023) have demonstrated remarkable generalization capabilities across diverse tasks (Raiaan et al. 2024), largely attributed to their ever-increasing scale. Adapting these powerful models to specific downstream applications is commonly achieved through fine-tuning (Han et al. 2024). However, conventional Full Fine-Tuning (FFT) approaches update all model parameters, making them computationally expensive and memory-intensive, thereby posing significant challenges for resource-limited deployments. As a solution, Parameter-Efficient Fine-Tuning (PEFT) (Pfeiffer et al. 2020) has been proposed to keep the majority of the pre-trained model’s parameters frozen and adjust only a small subset of new or existing parameters. Among various PEFT methods such as Adapters (Houlsby et al. 2019) and Prefix-Tuning (Li and Liang 2021), Low-Rank Adaptation (LoRA) (Hu et al. 2022) has gained particular prominence.
@@
LoRA, a simple yet effective PEFT method, hypothesizes that the changes in weight during adaptation tend to be low-rank, which is implemented by injecting trainable low-rank matrices into transformer layers. However, LoRA (shown in Fig. 1(a)) faces a fundamental low-rank bottleneck that consistently leads to performance gaps compared to full fine-tuning, particularly on complex reasoning tasks (Biderman et al. 2024). This limitation stems from the assumption of low-rank updating weights, which may not hold for tasks requiring substantial representation (Mao et al. 2025). When practitioners encounter insufficient performance, enlarging the rank number remains an alternative solution, but this directly increases the parameter and computational overhead, potentially compromising LoRA’s efficiency advantages.
@@
To mitigate LoRA’s limitations, the research community has explored various LoRA variants. Initially, researchers pursued capacity expansion through the Mixture-of-Experts technique (e.g., LoRAMoE) (Dou et al. 2024) (shown in Fig. 1(b)), which employs multiple parallel LoRA modules to decouple multi-task information. Although this enhances representational diversity by allowing different experts to specialize in distinct tasks, the independence of these experts limits the capture of shared knowledge and incurs substantial computation and parameter overhead. Subsequently, asymmetric architectures (Tian et al. 2024; Yang et al. 2025) emerged as a more parameter-efficient solution, exemplified by HydraLoRA’s “one-$A$, multi-BB” architecture, as shown in Fig. 1(c). This design is motivated by the observation that the down-projection matrix ($A$) tends to learn shared, general features by adapting to diverse data, while the up-projection matrix ($B$) captures task-specific variations. This paradigm of allocating representational complexity to the matrix $B$ has become a prevailing design philosophy. However, such a “single-$A$, multi-BB” strategy may introduce a potential representation bottleneck by relying on a single, low-rank $A$-matrix for all feature extraction. It appears that a single down-projection imposes a ceiling on the mutual information between the input and the update, thereby compressing every example into at most rank $r$ orthogonal directions (in Theorem 1). Empirical results corroborate this phenomenon: as illustrated in Fig. 2, a single‑$A$ matrix maps examples from three disparate tasks into largely overlapping regions, whereas an ensemble of experts $A$ produces well‑separated clusters suitable for multi-task adaptation.
@@
The above analysis motivates us to overcome LoRA’s representational bottleneck, requiring multiple down-projection matrices ($A$-experts) to capture diverse shared features, as shown in Fig. 1 (d). Meanwhile, the up-projection $B$ is used to integrate this knowledge for multi-domain and multi-task adaptations. To maintain architectural simplicity and computational efficiency, we employ only one up-projection matrix, which is demonstrated to be sufficient for effective adaptation while significantly reducing parameter overhead in experiments (Tab. 4). Motivated by the observations of feature similarity across adjacent transformer layers (Xiao et al. 2019) and further corroborated by our adjacent-layer CKA analysis (Fig. 3), we find that the outputs of $A$-experts are highly redundant across layers, indicating that these experts can be shared across layer groups. To this end, we consider designing a parameter-efficient sharing mechanism, which not only improves adaptation efficiency but also enables the model to learn more generalized features.
@@
Related Work
LoRA Architecture and Its Variants
@@
Parameter-Efficient Fine-Tuning (PEFT) has become essential for adapting Large Language Models (LLMs) to downstream tasks. Low-Rank Adaptation (LoRA) (Hu et al. 2022) only adds and updates one trainable pair of low-rank matrices $A$ and $B$ in each layer during fine-tuning, while keeping the existing weights frozen. It has been one of the most effective solutions in PEFT. However, the monolithic nature of a single LoRA module, with its simple one-to-one connection, limits its capacity to capture the diverse features required for complex tasks. This limitation spurred the development of Multi-LoRA architectures presented in Fig. 1. An early attempt, LoRAMoE (Dou et al. 2024), employs a Mixture-of-Experts framework with multiple LoRA modules. Although this enhances representational diversity, the independence of its experts impedes the learning of shared, domain-general knowledge. Alternatively, subsequent research shifted towards asymmetric structures that share certain components. Works like HydraLoRA (Tian et al. 2024) and MTL-LoRA (Yang et al. 2025) adopted a “one-$A$, multi-BB” paradigm, where a shared matrix $A$ captures common features and multiple matrices $B$ learn task-specific projections. Even more flexible structures like CoLA (Zhou, Yao, and Chen 2025), which allow for many-to-many relationships, still predominantly adhere to a “few-$A$, many-BB” philosophy. This places the primary burden of feature extraction on a smaller number of $A$-matrices, which can limit the model’s ability to capture the diverse and nuanced signals required for complex tasks. Our work explores an alternative approach by shifting the architecture to focus on the feature extraction stage. We employ a “multi-$A$, single-BB” structure, utilizing an ensemble of $A$-matrices to enrich the feature representations.
Cross-Layer Parameter Sharing Strategies
@@
Parameter sharing across layers has emerged as a fundamental strategy for reducing memory footprint and computational overhead while maintaining performance. ALBERT (Lan et al. 2019) achieved up to 90% parameter reduction in BERT by reusing attention and feed-forward weights across layers, while the Universal Transformer (Dehghani et al. 2018) formalized depth as recurrent time steps with shared parameters. Similar successes in computer vision (LaViT (Zhang et al. 2024)) and inference optimization (KVSharer (Yang et al. 2024)) demonstrate reductions of up to 4× in parameters and 30% in memory usage, respectively. Building upon these insights, the LoRA community has explored parameter sharing strategies tailored to low-rank adaptation. As models scale, the cumulative parameter count across layers has motivated sharing investigations. Recent approaches include VeRA’s (Kopiczko, Blankevoort, and Asano 2023) sharing of random matrices, VB-LoRA’s (Li, Han, and Ji 2024) vector-bank reconstruction, and BSLoRA’s (Zhou et al. 2025) granular slicing schemes. However, these methods face a fundamental limitation: their strategies assume symmetric LoRA structures, treating the $A$ and $B$ matrices equivalently. This prevents them from exploiting the functional asymmetry between feature extraction and projection identified. Guided by empirical analysis of inter-layer LoRA similarity, our method shares the down-projection matrices ($A$) across layer groups while preserving layer-specific up-projection matrices ($B$).
@@
The Proposed Method
Preliminaries
Low-Rank Adaptation (LoRA) (Hu et al. 2022) posits that the additional and updated weight $\Delta W$ is low-rank for a pre-trained weight matrix $W\in\mathbb{R}^{d_{out}\times d_{in}}$ during fine-tuning. Specifically, $\Delta W$ is decomposed into the low-rank matrices, $A\in\mathbb{R}^{r\times d_{in}}$ and $B\in\mathbb{R}^{d_{out}\times r}$, where the rank $r\ll\min(d_{in},d_{out})$. During adaptation, $W$ remains frozen, and only $A$ and $B$ are trained. The modified forward pass for a hidden state $x$ is given by:
@@
For initialization, the matrix $A$ is typically initialized with the Kaiming Uniform (He et al. 2015), while $B$ is initialized to zero, ensuring that $\Delta W=BA$ is zero at the beginning of training. Functionally, the LoRA module can be conceptualized as a two-stage process: feature compression and decompression. The matrix $A$ acts as a projector or compressor, mapping the high-dimensional input features into a low-rank, compact space. Subsequently, the matrix $B$ acts as a restorer or decompressor, projecting these low-rank features back to the original output dimension space. For stable training, the LoRA update is typically scaled by a constant $\frac{\alpha}{r}$, where $\alpha$ is a learning rate-like scalar. Thus, the final forward pass is $h=W_{0}x+\frac{\alpha}{r}BAx$.
Representational Bottleneck of Single-$A$ LoRA
Theorem 1 (Information Ceiling).
For any input ${\mathbf{x}}\!\in\!\mathbb{R}^{d_{\text{in}}}$ with covariance matrix $\Sigma_{x}\succeq 0$, the down‑projection ${\mathbf{u}}=A\mathbf{x}$ $(A\!\in\!\mathbb{R}^{r\times d_{\text{in}}})$ in a rank‑$r$ LoRA layer satisfies:
@@
$I(\mathbf{x};B\mathbf{u})\leq\tfrac{1}{2}\log\det\!\bigl(I_{r}+A\Sigma_{x}A^{\top}\bigr),\newline \text{rank}(\mathbf{u})\leq r,$ (2)
for any deterministic post‑mapping $B$ (even multiple BB‑heads). Thus, a single‑$A$ LoRA can transmit at most $r$ orthogonal information channels.
Proof (sketch).
Because g=Bg\!=\!B is deterministic, ${\mathbf{x}}\!\to\!{\mathbf{u}}\!\to\!g({\mathbf{u}})$ forms a Markov chain; The data‑processing inequality yields the information bound. The rank bound follows from $\text{rank}(A)=r$. The full proof is provided in Appendix A. ∎
Remark. Replacing $A$ by an ensemble $\{A_{i}\}_{i=1}^{N}$ and aggregating $\tilde{\mathbf{u}}=\sum_{i}A_{i}\mathbf{x}$ preserves the rank constraint, but increases $\det(I_{r}+\!\sum_{i}\!A_{i}\Sigma_{x}A_{i}^{\top})$, thereby lifting the ceiling in (2). This insight motivates our “multi‑$A$, single‑BB” design, which diversifies the feature extractor without inflating projection heads.
The Multi-$A$ Expert (MAE) Block

### v2

@@ Abstract
Low-Rank Adaptation (LoRA) has emerged as a dominant method in Parameter-Efficient Fine-Tuning (PEFT) for large language models, which augments the transformer layer with one down-projection $A$ and one up-projection $B$. However, LoRA’s reliance on a single down-projection matrix ($A$) creates a representational bottleneck, as this solitary feature extractor is inherently insufficient for capturing the diverse signals required by complex tasks. This motivates our architectural shift to focus on enriching the feature adaptation to improve the downstream task adaptation ability. We propose MASA (Multi-$A$ Shared Adaptation), an architecture that implements a “multi-$A$, single-BB” structure where the multi-$A$ expert ensemble is asymmetrically shared across layers to ensure parameter efficiency. In MASA, these specialized experts capture diverse features, which are then integrated by a single, layer-specific $B$-matrix. The effectiveness and versatility of our method are validated through a comprehensive suite of experiments spanning multi-domain generalization, single-domain specialization, and multi-task reasoning. For example, on the MMLU benchmark, MASA achieves an average accuracy of 59.62%, outperforming the standard LoRA by 1.08 points (a relative improvement of 1.84%) while training only 0.52% of the parameters.
1 Introduction
@@
Figure 1: Architectures of LoRA variants and our proposed MASA. (a) LoRA: Each fine-tuned layer is augmented with a single pair of low‑rank adapters, one down‑projection AA and one up‑projection $B$. (b) LoRA+MoE: Model capacity is increased by instantiating $k$ independent adapter pairs $(A_{i},B_{i})$. (c) HydraLoRA: Several up‑projection heads $B_{i}$ share a common down‑projection AA, forming a “single‑AA, multi‑BB” topology for parameter reuse. (d) Ours: We employ one shared up‑projection $B$ with multiple down‑projections $A_{i}$, which offers a balanced trade‑off between efficiency and representational capacity.
@@
Large Language Models (LLMs) (Brown et al., 2020; Devlin et al., 2019; Touvron et al., 2023; OpenAI, 2023) have demonstrated remarkable generalization capabilities across diverse tasks (Raiaan et al., 2024), largely attributed to their ever-increasing scale. Adapting these powerful models to specific downstream applications is commonly achieved through fine-tuning (Han et al., 2024). However, conventional Full Fine-Tuning (FFT) approaches update all model parameters, making them computationally expensive and memory-intensive, thereby posing significant challenges for resource-limited deployments. As a solution, Parameter-Efficient Fine-Tuning (PEFT) (Pfeiffer et al., 2020) has been proposed to keep the majority of the pre-trained model’s parameters frozen and adjust only a small subset of new or existing parameters. Among various PEFT methods such as Adapters (Houlsby et al., 2019) and Prefix-Tuning (Li and Liang, 2021), Low-Rank Adaptation (LoRA) (Hu et al., 2022) has gained particular prominence.
@@
LoRA, a simple yet effective PEFT method, hypothesizes that the changes in weight during adaptation tend to be low-rank, which is implemented by injecting trainable low-rank matrices into transformer layers. However, LoRA (shown in Fig. 1(a)) faces a fundamental low-rank bottleneck that consistently leads to performance gaps compared to full fine-tuning, particularly on complex reasoning tasks (Biderman et al., 2024). This limitation stems from the assumption of low-rank updating weights, which may not hold for tasks requiring substantial representation (Mao et al., 2025). When practitioners encounter insufficient performance, increasing the rank remains an alternative solution, but this directly increases the parameter and computational overhead, potentially compromising LoRA’s efficiency advantages.
@@
To mitigate LoRA’s limitations, the research community has explored various LoRA variants. Initially, researchers pursued capacity expansion through the Mixture-of-Experts technique (e.g., LoRAMoE) (Dou et al., 2024) (shown in Fig. 1(b)), which employs multiple parallel LoRA modules to decouple multi-task information. Although this enhances representational diversity by allowing different experts to specialize in distinct tasks, the independence of these experts limits the capture of shared knowledge and incurs substantial computation and parameter overhead. Subsequently, asymmetric architectures (Tian et al., 2024; Yang et al., 2025) emerged as a more parameter-efficient solution, exemplified by HydraLoRA’s “one-$A$, multi-BB” architecture, as shown in Fig. 1(c). This design is motivated by the observation that the down-projection matrix ($A$) tends to learn shared, general features by adapting to diverse data, while the up-projection matrix ($B$) captures task-specific variations. This paradigm of allocating representational complexity to the matrix $B$ has become a prevailing design philosophy. However, such a “single-$A$, multi-BB” strategy may introduce a potential representation bottleneck by relying on a single, low-rank $A$-matrix for all feature extraction. It appears that a single down-projection imposes a ceiling on the mutual information between the input and the update, thereby compressing every example into at most rank $r$ orthogonal directions (in Theorem 3.1). Empirical results corroborate this phenomenon via t-SNE (Maaten and Hinton, 2008): As illustrated in Fig. 2, a single‑$A$ matrix maps examples from three disparate tasks into largely overlapping regions, whereas an ensemble of experts $A$ produces well‑separated clusters suitable for multi-task adaptation.
@@
The above analysis motivates us to overcome LoRA’s representational bottleneck, requiring multiple down-projection matrices ($A$-experts) to capture diverse shared features, as shown in Fig. 1 (d). Meanwhile, the up-projection $B$ is used to integrate this knowledge for multi-domain and multi-task adaptations. To maintain architectural simplicity and computational efficiency, we employ only one up-projection matrix, which is demonstrated to be sufficient for effective adaptation while significantly reducing parameter overhead. Motivated by the observations of feature similarity across adjacent transformer layers (Xiao et al., 2019) and further corroborated by our adjacent-layer CKA (Kornblith et al., 2019), we find that the outputs of $A$-experts are highly redundant across layers, indicating that these experts can be shared across layer groups (Fig. 3). To this end, we consider designing a parameter-efficient sharing mechanism, which not only improves adaptation efficiency but also enables the model to learn more generalized features.
@@
2 Related Work
2.1 LoRA Architecture and Its Variants
@@
Parameter-Efficient Fine-Tuning (PEFT) has become essential for adapting Large Language Models (LLMs) to downstream tasks. Low-Rank Adaptation (LoRA) (Hu et al., 2022) only adds and updates one trainable pair of low-rank matrices $A$ and $B$ in each layer during fine-tuning, while keeping the existing weights frozen. It has been one of the most effective solutions in PEFT. However, the monolithic nature of a single LoRA module, with its simple one-to-one connection, limits its capacity to capture the diverse features required for complex tasks. This limitation spurred the development of Multi-LoRA architectures presented in Fig. 1. An early attempt, LoRAMoE (Dou et al., 2024), employs a Mixture-of-Experts framework with multiple LoRA modules. Although this enhances representational diversity, the independence of its experts impedes the learning of shared, domain-general knowledge. Alternatively, subsequent research shifted towards asymmetric structures that share certain components. Works like HydraLoRA (Tian et al., 2024) and MTL-LoRA (Yang et al., 2025) adopt a “one-$A$, multi-BB” paradigm, where a shared matrix $A$ captures common features and multiple matrices $B$ learn task-specific projections. Even more flexible structures like CoLA (Zhou et al., 2025a), which allow for many-to-many relationships, still predominantly adhere to a “few-$A$, many-BB” philosophy. This places the primary burden of feature extraction on a smaller number of $A$-matrices, which can limit the model’s ability to capture the diverse and nuanced signals required for complex tasks. Our work explores an alternative approach by shifting the architecture to focus on the feature extraction stage. We employ a “multi-$A$, single-BB” structure, utilizing an ensemble of $A$-matrices to enrich the feature representations.
2.2 Cross-Layer Parameter Sharing Strategies
@@
Parameter sharing across layers has emerged as a fundamental strategy for reducing memory footprint and computational overhead while maintaining performance. ALBERT (Lan et al., 2019) achieved up to 90% parameter reduction in BERT by reusing attention and feed-forward weights across layers, while the Universal Transformer (Dehghani et al., 2018) formalized depth as recurrent time steps with shared parameters. Similar successes in computer vision (LaViT (Zhang et al., 2024)) and inference optimization (KVSharer (Yang et al., 2024)) demonstrate reductions of up to 4× in parameters and 30% in memory usage, respectively. Building upon these insights, the LoRA community has explored parameter sharing strategies tailored to low-rank adaptation. As models scale, the cumulative parameter count across layers has motivated sharing investigations. Recent approaches include VeRA’s (Kopiczko et al., 2023) sharing of random matrices, VB-LoRA’s (Li et al., 2024) vector-bank reconstruction, and BSLoRA’s (Zhou et al., 2025b) granular slicing schemes. However, these methods face a fundamental limitation: their strategies assume symmetric LoRA structures, treating the $A$ and $B$ matrices equivalently. This prevents them from exploiting the functional asymmetry between feature extraction and projection identified. Guided by empirical analysis of inter-layer LoRA similarity, our method shares the down-projection matrices ($A$) across layer groups while preserving layer-specific up-projection matrices ($B$).
@@
3 The Proposed Method
3.1 Preliminaries
Low-Rank Adaptation (LoRA) (Hu et al., 2022) posits that the additional and updated weight $\Delta W$ is low-rank for a pre-trained weight matrix $W\in\mathbb{R}^{d_{out}\times d_{in}}$ during fine-tuning. Specifically, $\Delta W$ is decomposed into the low-rank matrices, $A\in\mathbb{R}^{r\times d_{in}}$ and $B\in\mathbb{R}^{d_{out}\times r}$, where the rank $r\ll\min(d_{in},d_{out})$. During adaptation, $W$ remains frozen, and only $A$ and $B$ are trained. The modified forward pass for a hidden state $x$ is given by:
@@ 3.1 Preliminaries
For initialization, the matrix $A$ is typically initialized with the Kaiming Uniform (He et al., 2015), while $B$ is initialized to zero, ensuring that $\Delta W=BA$ is zero at the beginning of training. Functionally, the LoRA module can be conceptualized as a two-stage process: feature compression and decompression. The matrix $A$ acts as a projector or compressor, mapping the high-dimensional input features into a low-rank, compact space. Subsequently, the matrix $B$ acts as a restorer or decompressor, projecting these low-rank features back to the original output dimension space. For stable training, the LoRA update is typically scaled by a constant $\frac{\alpha}{r}$, where $\alpha$ is a learning rate-like scalar. Thus, the final forward pass is $h=W_{0}x+\frac{\alpha}{r}BAx$.
3.2 A Rigorous Bottleneck of Single-Extractor LoRA
LoRA updates a frozen linear map via $\Delta W=BA$, where $A\in\mathbb{R}^{r\times d_{\text{in}}}$ acts as an extractor producing $\mathbf{u}=A\mathbf{x}\in\mathbb{R}^{r}$, and $B$ (or any deterministic $g$) only remixes these $r$ features. Hence the adaptation branch is constrained by an $r$-dimensional bottleneck shaped by $A$.
Finite precision / noisy representation. For continuous variables, mutual information under deterministic mappings may be ill-defined or infinite. We therefore adopt a standard finite-precision/noisy-representation view:
$\tilde{\mathbf{u}}=\mathbf{u}+\boldsymbol{\varepsilon},\qquad\boldsymbol{\varepsilon}\sim\mathcal{N}(\mathbf{0},\sigma^{2}I_{r}),\ \boldsymbol{\varepsilon}\perp\mathbf{x}.$ (2)
Theorem 3.1 (Information Ceiling (finite precision)).
Let $\mathbf{x}\in\mathbb{R}^{d_{\text{in}}}$ have covariance $\Sigma_{x}\succeq 0$ and let $\mathbf{u}=A\mathbf{x}$ with $A\in\mathbb{R}^{r\times d_{\text{in}}}$. Define $\tilde{\mathbf{u}}$ as in (2). Then for any deterministic post-mapping $g$ (including $g(\tilde{\mathbf{u}})=B\tilde{\mathbf{u}}$ and multiple $B$-heads),
@@ 3.2 A Rigorous Bottleneck of Single-Extractor LoRA
$I(\mathbf{x};g(\tilde{\mathbf{u}}))\;\leq\;\tfrac{1}{2}\log\det\!\Bigl(I_{r}+\tfrac{1}{\sigma^{2}}A\Sigma_{x}A^{\top}\Bigr),\quad\mathrm{rank}(\mathbf{u})\leq r.$ (3)
Proof.
[…]
3.3 The Whole MASA Framework
Based on Theorem 3.1, we find that the single matrix $A$ in standard LoRA and its asymmetric variants (e.g., HydraLoRA) may become a representational bottleneck, limiting the model’s capacity to extract diverse and nuanced features. Moreover, the across-layer similarity on the feature extractor $A$ often incurs redundant computation. To balance performance and efficiency, we design the Multi-$A$ Expert block with our asymmetric sharing strategy to construct Multi-$A$ Shared Adaptation (MASA) framework. The overall architecture is illustrated in Fig. 4, which consists of two key modules, Multi-$A$ Expert Block (MAE) and Asymmetric Cross-layer Sharing (ACS). The shared matrices $A_{i}$ are managed by MAE, enabling inter-layer sharing to improve parameter efficiency via ACS. In detail, MAE and ACS are described in Section 3.4 and  3.5.
$\small h^{(l)}=W_{0}^{(l)}x+\frac{\alpha}{r}B^{(l)}\left(\sum_{i=1}^{N}A_{i}^{(k)}\right)x,\quad\text{where }k=\lfloor l/S\rfloor.$ (4)
Here, $k$ is the index of the shared $A$-matrix group, determined by the layer $l$ and the size of sharing group $S$.
3.4 The Multi-$A$ Expert Block